A plane traveled 580 miles to Ankara and back. The trip there was with the wind. It took 5 hours. The trip back was into the wind. The trip took 10 hours. Find the speed of the plane in still air and the speed of the wind. i know this should be relatively easy, but for some reason I am having a brain block and can't seem to make any headway on this problem. Any help you can provide will be GREATLY appreciated!
Answer by scott8148(5879)

rate = distance / time

p + w = 580 / 5 = 116

p - w = 580 / 10 = 58

adding the equations ___ 2p = 174 ___ p = 87

substituting ___ 87 + w = 116 ___ w = 29